OpenStudy (anonymous):

x^2-8x+7 factor the trinomial

OpenStudy (jennychan12):

(x-7)(x-1) -7*-1 = 7 -7+(-1) = -8

OpenStudy (anonymous):

so is that it -8

OpenStudy (anonymous):

so jenychan is it -8

OpenStudy (jennychan12):

no. it's (x-7)(x-1) you're looking for two numbers that add up to -8 but multiply to 7 (if you want you can set up xn = 7 and x+n = -8) WORK: x*x = x^2 -7*-1 = 7 -7+(-1) = -8

OpenStudy (anonymous):

4 and 2

OpenStudy (jennychan12):

..... no. 4 and 2 don't add up to 8 nor do they multiply to 7 (7 is a prime number)

OpenStudy (anonymous):

maybe it cannot be factored

OpenStudy (anonymous):

i do not get it

OpenStudy (jennychan12):

so you have x^2-8x+7 right? the middle term (-8x) means that you're looking for two numbers that'll add up to -8. There are many possibilities. Don't look at this term first. Let's look at +7. What times what gives you 7? Well, we know that since 7 is a prime number, you can only have 7*1 or -7*-1 right? But 7+1 = +8, not -8, so our only choice is -7 and -1 so it'd be (x-7)(x-1)

OpenStudy (anonymous):

thanks